# Field-orientation-dependent magnetic phases in GdRu2Si2 probed with muon-spin spectroscopy

### v2

@@ Abstract
Centrosymmetric GdRu2Si2 exhibits a variety of multi-Q magnetic states as a function of temperature and applied magnetic field, including a square skyrmion-lattice phase. The material’s behavior is strongly dependent on the direction of the applied field, with different phase diagrams resulting for fields applied parallel or perpendicular to the crystallographic $c$ axis. Here, we present the results of muon-spin relaxation ($\mu^{+}$) measurements on single crystals of GdRu2Si2. Our analysis is based on the computation of muon stopping sites and consideration of zero-point motion effects, allowing direct comparison with the underlying spin textures in the material. Using transverse-field $\mu^{+}$ with fields applied along either the [001] or [100] crystallographic directions, we distinguish between the magnetic phases in this system via their distinct muon response, providing additional evidence for the skyrmion and meron-lattice phases, while also suggesting the existence of RKKY-driven muon hyperfine coupling. Zero-field $\mu^{+}$ provides clear evidence for a transition between two distinct magnetically-ordered phases at 39 K.
@@ I Introduction
Skyrmions are topologically-protected vortices of magnetization that exist in some magnetically-ordered materials [1]. The search for new materials that host a magnetic skyrmion state has been motivated by the numerous potential applications of skyrmions in the areas of ultra-efficient information storage and spintronics [2]. Skyrmions are typically stabilized by the Dzyaloshinskii–Moriya interaction (DMI), which can only occur in systems that lack inversion symmetry. Magnets based on Gd (possessing 4f7 core states with $S=7/2$, $L=0$) are promising candidates for realizing non-collinear spin structures such as skyrmions, since the moments show non-preferential orientation. Two Gd-based magnets where skyrmions have recently been discovered are Gd2PdSi3 [3] and Gd3Ru4Al12 [4]. In these centrosymmetric materials, skyrmions have been proposed to be stabilized by either short- or long-range magnetic frustration rather than DMI. This can lead to interesting competition with other magnetic phases, and potentially new behavior of the skyrmion state. Skyrmions have also been observed in GdRu2Si2 [5], another centrosymmetric Gd-based magnet. However, the structure of GdRu2Si2, shown in Fig. 1(a), comprises alternating square-lattice Gd and Ru2Si2 layers and therefore, unlike for Gd2PdSi3 and Gd3Ru4Al12, short-range geometrical frustration cannot be used to explain the formation of the skyrmion lattice (SkL). Magnetic resonant x-ray scattering (RXS) and real space using Lorentz transmission electron microscopy reveal the existence of a double-Q magnetic state, corresponding to a square SkL [5]. This is distinct from the hexagonal SkL found in Gd2PdSi3 and GdRu4Al12, and in the vast majority of the noncentrosymmetric skyrmion-hosting materials. Moreover, the existence of a square SkL has only been demonstrated [5] for applied fields $\bm{B}_{0}\parallel$ [001], for which the SkL phase is stabilized at $\bm{B}_{0}\approx$ 2 T. For fields in the $ab$ plane, a different phase diagram is obtained [6, 7]. The magnetic phase diagrams for $\bm{B}_{0}\parallel$ [001] and $\bm{B}_{0}\parallel$ [100] have been studied using RXS [7], with these measurements revealing a rich variety of double-Q magnetic structures, including the antiferroic order of meron/antimeron-like textures. Neutron diffraction measurements on single-crystal and polycrystalline 160GdRu2Si2 found, in addition to scattering at the principal propagation vectors $\bm{q}_{1}$ and $\bm{q}_{2}$, higher-order satellites at $\bm{q}_{1}+2\bm{q}_{2}$, which prompted a revision of the zero-field magnetic ground state to a double-Q constant-moment magnetic structure, with one-dimensional topological charge density [8]. At small values of applied field, an additional phase has been identified just below $T_{\mathrm{N}}$ whose properties have not been reported.
@@
00footnotetext: See Supplemental Material at [URL will be inserted by publisher] for further details of the the density functional theory calculations and simulations of the magnetic field distributions, which includes Refs. [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42].
@@
Here, we use muon-spin spectroscopy ($\mu^{+}$) [9] to probe the magnetic phases in GdRu2Si2 as a function of magnetic field and temperature. The $\mu^{+}$ technique has been shown to be sensitive to the distribution of magnetic fields arising from the SkL phase in Cu2OSeO3 [10], as well as those found in the helical [11] and conical [12] phases of skyrmion-hosting MnSi. Implanted muons have also been shown to effectively probe the emergent dynamics associated with Bloch SkL, such as in Cu2OSeO3 and CoxZnyMn20-x-y [13] and in Gd2PdSi3 [14], or those accompanying a Néel SkL, as found in GaV4S8 and GaV4Se8 [15, 16]. Using transverse-field (TF) $\mu^{+}$ for magnetic fields applied along either [001] or [100], we show here how the response of the implanted muon reflects the occurrence of skyrmion and other double-Q magnetic phases. Zero-field (ZF) $\mu^{+}$ measurements provide clear evidence for a transition between two distinct magnetically-ordered phases at 39 K, showing that the phase reported just below $T_{\mathrm{N}}=45$ K is ordered throughout the bulk of the material. Our modelling also affords an estimate of the low-temperature ordered moment of $4.8(1)\mu_{\mathrm{B}}$ and shows the influence of long-range Ruderman-Kittel-Kasuya-Yosida (RKKY) coupling in the system, via the muon’s coupling to the electronic degrees of freedom.
@@ II Muon site calculations
Implanted muons are sensitive to the distribution of magnetic fields at the muon sites. Knowledge of these sites can facilitate a quantitative understanding of $\mu^{+}$ measurements. Progress has recently been made in computing stopping sites using density DFT [18], although treatment of quantum-mechanical effects such as zero-point motion (ZPM) has hitherto been limited to special cases [19, 9]. We have used DFT methods to calculate the muon sites in GdRu2Si2 [17] and find two crystallographically-distinct candidate sites [Fig. 1(a)]. The calculated entanglement witnesses [19] at each of these sites places both of these muon states firmly within the weakly-bound adiabatic limit. However, on taking into account the quantum zero-point energy, a muon at site 2 (the higher-energy state of the two) has sufficient energy to overcome the energy barrier between its classical location and that of site 1 [Fig. 1(c)]. Site 2 is therefore not stable under quantum fluctuations of the muon position. While the barrier from site 1 to site 2 is sufficiently high that we do not expect muons at site 1, we also need to consider the possiblity of quantum mechanical hopping. By considering the effective one-dimensional potential for the muon along the minimum-energy path between the classical turning points, i.e., where the black and red lines in Fig. 1(c) intersect, we used the WKB approximation to estimate a tunnelling rate $\approx$0.05 Hz for muons at site 1 [17], which is significantly longer than the timescale over which measurements are made. We therefore conclude that there is a single crystallographically-distinct stable stopping site in this material and we plot the displacements of the nearby ions in the vicinity of this site in Fig. 1(b). The muon does not introduce any significant perturbations to its local environment, with all displacements $<0.1$, and is therefore expected to be a faithful probe of the magnetism in this system. These calculations do not take into account the possible effect of charge density waves, which have been proposed to result from the coupling between itinerant electrons and localized moments in this system [20], and could result in small perturbations to the muon site.
@@
Figure 1: (a) Two crystallographically distinct muon stopping sites in GdRu2Si2, obtained using DFT. (b) Muon-induced displacements of ions as a function of distance from the lowest-energy muon site. (c) Energy barriers between muon sites. Red and blue lines indicate the total (classical plus quantum, zero-point) energy of the muon at sites 1 and 2, respectively.
@@ III Results and Discussion
III.1 TF $\bm{\mu}^{+}$ measurements
@@
Transverse-field $\mu^{+}$ measurements were made on aligned crystals of GdRu2Si2 after zero-field cooling. In these measurements, the muon spin precesses about the sum of the external magnetic field, which is oriented perpendicular to the initial muon-spin direction, and any internal magnetic fields present in the sample. Two sets of measurements were made: one with the crystals oriented such that applied field $\bm{B}_{0}$ pointed along [001] and another set with the crystals oriented such that $\bm{B}_{0}$ pointed along the [100] direction.
@@
We first consider the spectra measured at $T=40$ K, where we can access the field-polarized state. Fourier transform (FT) $\mu^{+}$ spectra measured at different fields are shown in Fig. 2 for each of the orientations. For $\bm{B}_{0}\parallel$ [001] the spectra comprise a peak close to the applied field and a secondary peak at lower fields which becomes more prominent as the field is increased. (We note that the central peak is split in two for all of the TF spectra measured on this material, which could result from a fraction of the muons stopping in parts of the sample where the internal field almost, but not completely, cancels, such as, for example, domain walls or surfaces.) For $\bm{B}_{0}\parallel$ [100] there is an additional peak at fields larger than $B_{0}$ that is absent for the other orientation. We therefore fit the spectra for fields along [001] and [100] to the sum of two or three Lorentzians, respectively, and plot the behavior of the peak centers in Figs. 2(d) and 2(h). We see that the magnitudes of the field shifts at the two peaks initially increases approximately linearly with the applied field, but then begins to level off, suggesting that the magnetization is starting to saturate. We also note the absence of any satellite peaks for $B_{0}=0.5$ T $\parallel$ [001] in Fig. 2(a), with the magnetic phase diagram for this orientation [shown in Fig. 3(i)] having a phase boundary between 0.5 and 1 T at this temperature. This suggests that the low-field, higher-temperature magnetically ordered state gives rise to a broad distribution of magnetic fields at the muon site, relaxing the contribution of muons stopping in the sample out of the TF spectra, a scenario which is consistent with our ZF spectra, which show a large relaxation rate at this temperature. On the other hand, spectra measured at 0.5 and 1.0 T for $\bm{B}_{0}\parallel[100]$, shown in Figs. 2(e) and 2(f), respectively, are expected to occupy the same magnetic phase [as seen in our ac susceptibility measurements shown in Fig. 2(j)] and both show satellite peaks. This suggests the zero-field ordered state survives up to higher fields applied along [001] than for those along [100].
@@
The number of peaks observed for each of the field orientations reflects the symmetry of the muon stopping site. The local magnetic field at the muon site is the sum of the externally applied field and dipolar, Lorentz, demagnetizing and hyperfine fields. For Gd moments along $c$, muons at site 1 will all experience an internal field of $(-0.063-0.140N+7.77A)\mu_{\mathrm{Gd}}$ T along $c$, where the first term is sum of the dipolar and Lorentz fields, the second term is the demagnetizing field, with the demagnetizing factor $N$ depending on the sample shape, and the final term is the hyperfine field, which is proportional to the hyperfine coupling constant $A$, here measured in Å-3. All of these contributions to the internal field are proportional to the Gd magnetic moment $\mu_{\mathrm{Gd}}$, which in this expression is measured in Bohr magnetons. On the other hand, for moments along $a$, instances of site 1 lying on the $b$ axis will experience a field $(0.295-0.140N+7.77A)\mu_{\mathrm{Gd}}$ T along $a$, whereas the crystallographically equivalent site lying on the $a$ axis will instead experience a field $(-0.091-0.140N+7.77A)\mu_{\mathrm{Gd}}$ T. For $\bm{B}_{0}\parallel$ [100] the separation of the satellite peaks, $\Delta B$, is therefore related to $\mu_{\mathrm{Gd}}$ by $\Delta B=0.386\mu_{\mathrm{Gd}}$ T. The development of $\mu_{\mathrm{Gd}}$ as a function of applied magnetic field is shown in Fig. 2(i). We have also measured the Gd moment as a function of magnetic field using magnetometry [17] [Fig. 2(i)]. The values of $\mu_{\mathrm{Gd}}$ obtained using these approaches show reasonable agreement, though we note that those obtained from magnetometry are consistently larger and have a slightly different functional dependence with applied field. Taking the sample geometry to be a flat plate with a perpendicular magnetization, we assume $N\approx 1$. Fitting either the positive or negative and field shifts for $\bm{B}_{0}\parallel$ [100] to a straight line in $\mu_{\mathrm{Gd}}(B_{0})$ [blue or red lines in Fig. 2(h), respectively] yields $A=-0.004~{}\mathrm{\AA}^{-3}$. This negative hyperfine contribution can be explained by the spin-polarization of conduction electrons through the RKKY interaction, which has been predicted to play an important role in stabilizing the skyrmion lattice in GdRu2Si2 [21]. We note that while the multiplicity of the peaks for fields along [100] is not specific to the candidate sites found in our calculations, it does rule out sites with axial symmetry, such as the site at (0.5, 0.5, 0) that has been proposed for the isostructural compound CeRu2Si2 [22], which would exhibit only a single peak.

### v3

@@ Abstract
Centrosymmetric GdRu2Si2 exhibits a variety of multi-Q magnetic states as a function of temperature and applied magnetic field, including a square skyrmion-lattice phase. The material’s behavior is strongly dependent on the direction of the applied field, with different phase diagrams resulting for fields applied parallel or perpendicular to the crystallographic $c$ axis. Here, we present the results of muon-spin relaxation ($\mu^{+}$) measurements on single crystals of GdRu2Si2. Our analysis is based on the computation of muon stopping sites and consideration of zero-point motion effects, allowing direct comparison with the underlying spin textures in the material. The muon site is confirmed experimentally, using angle-dependent measurements of the muon Knight shift. Using transverse-field $\mu^{+}$ with fields applied along either the [001] or [100] crystallographic directions, we distinguish between the magnetic phases in this system via their distinct muon response, providing additional evidence for the skyrmion and meron-lattice phases, while also suggesting the existence of RKKY-driven muon hyperfine coupling. Zero-field $\mu^{+}$ provides clear evidence for a transition between two distinct magnetically-ordered phases at 39 K.
@@ I Introduction
Skyrmions are topologically-protected vortices of magnetization that exist in some magnetically-ordered materials [1]. The search for new materials that host a magnetic skyrmion state has been motivated by the numerous potential applications of skyrmions in the areas of ultra-efficient information storage and spintronics [2]. Skyrmions are typically stabilized by the Dzyaloshinskii–Moriya interaction (DMI), which can only occur in systems that lack inversion symmetry. Magnets based on Gd (possessing 4f7 core states with $S=7/2$, $L=0$) are promising candidates for realizing non-collinear spin structures such as skyrmions, since the moments show non-preferential orientation. Two Gd-based magnets where skyrmions have recently been discovered are Gd2PdSi3 [3] and Gd3Ru4Al12 [4]. In these centrosymmetric materials, skyrmions have been proposed to be stabilized by either short- or long-range magnetic frustration rather than DMI. This can lead to interesting competition with other magnetic phases, and potentially new behavior of the skyrmion state. Skyrmions have also been observed in GdRu2Si2 [5], another centrosymmetric Gd-based magnet. However, the structure of GdRu2Si2, shown in Fig. 1(a), comprises alternating square-lattice Gd and Ru2Si2 layers and therefore, unlike for Gd2PdSi3 and Gd3Ru4Al12, short-range geometrical frustration cannot be used to explain the formation of the skyrmion lattice (SkL). Magnetic resonant x-ray scattering (RXS) and Lorentz transmission electron microscopy reveal the existence of a double-Q magnetic state, corresponding to a square SkL [5]. This is distinct from the hexagonal SkL found in Gd2PdSi3 and GdRu4Al12, and in the vast majority of the noncentrosymmetric skyrmion-hosting materials. Moreover, the existence of a square SkL has only been demonstrated [5] for applied fields $\bm{B}_{0}\parallel$ [001], for which the SkL phase is stabilized at $B_{0}\approx$ 2 T. For fields in the $ab$ plane, a different phase diagram is obtained [6, 7]. The magnetic phase diagrams for $\bm{B}_{0}\parallel$ [001] and $\bm{B}_{0}\parallel$ [100] have been studied using RXS [7], with these measurements revealing a rich variety of double-Q magnetic structures, including the antiferroic order of meron/antimeron-like textures. Neutron diffraction measurements on single-crystal and polycrystalline 160GdRu2Si2 found, in addition to scattering at the principal propagation vectors $\bm{q}_{1}$ and $\bm{q}_{2}$, higher-order satellites at $\bm{q}_{1}+2\bm{q}_{2}$. This prompted a revision of the zero-field magnetic ground state to a double-Q constant-moment magnetic structure, with one-dimensional topological charge density [8]. At small values of applied field, an additional phase has been identified just below $T_{\mathrm{N}}$ whose properties are less well understood, but which was recently revealed to involve a single magnetic propagation vector that splits into two wave vectors with different magnitudes upon further cooling [9].
@@
00footnotetext: See Supplemental Material at [URL will be inserted by publisher] for further details of the the density functional theory calculations and simulations of the magnetic field distributions, which includes Refs. [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 21, 42, 43, 44].
@@
Here, we use muon-spin spectroscopy ($\mu^{+}$) [10] to probe the magnetic phases in GdRu2Si2 as a function of magnetic field and temperature. The $\mu^{+}$ technique has been shown to be sensitive to the distribution of magnetic fields arising from the SkL phase in Cu2OSeO3 [11], as well as those found in the helical [12] and conical [13] phases of skyrmion-hosting MnSi. Implanted muons have also been shown to effectively probe the emergent dynamics associated with Bloch SkL, such as in Cu2OSeO3 and CoxZnyMn20-x-y [14] and in Gd2PdSi3 [15], or those accompanying a Néel SkL, as found in GaV4S8 and GaV4Se8 [16, 17]. Using transverse-field (TF) $\mu^{+}$ for magnetic fields applied along either [001] or [100], we show here how the response of the implanted muon reflects the occurrence of skyrmion and other double-Q magnetic phases. Zero-field (ZF) $\mu^{+}$ measurements provide clear evidence for a transition between two distinct magnetically-ordered phases at 39 K, showing that the phase reported just below $T_{\mathrm{N}}=45$ K is ordered throughout the bulk of the material. Our modelling, which includes muon-site determination and its confirmation via experiment, also provides an estimate of the low-temperature ordered moment of $4.8(1)\mu_{\mathrm{B}}$ and shows the influence of long-range Ruderman-Kittel-Kasuya-Yosida (RKKY) coupling in the system, via the muon’s coupling to the electronic degrees of freedom.
@@ II Muon site calculations
Implanted muons are sensitive to the distribution of magnetic fields at the muon stopping sites. Knowledge of these sites can facilitate a quantitative understanding of $\mu^{+}$ measurements. Progress has recently been made in computing stopping sites using DFT [19], although treatment of quantum-mechanical effects such as zero-point motion (ZPM) has hitherto been limited to special cases [20, 10]. We have used DFT methods to calculate the muon sites in GdRu2Si2 [18] and find two crystallographically-distinct candidate sites [Fig. 1(a)]. These sites were further investigated to evaluate their stability. In the weakly-bound adiabatic limit (i.e., with zero entanglement with nearby nuclear positions) [20], the zero-point energy (ZPE) of the muon can be approximated as the sum of the three highest frequency phonon modes. The entanglement witness $w_{1}$, obtained by projecting the three highest-energy (normalized) phonon normal modes onto pure muon motion, summing the squared norms of the resulting (projected) phonon eigenvectors, and subtracting the value 3 [20], can be used to determine whether the muon is in fact in this limit. This results in $w=0$ for zero muon–nuclear entanglement (weakly-bound adiabatic limit) and $w_{1}<0$ in the entangled, many-body case. We obtained entanglement witnesses $w_{1}=-0.0079$ and $w_{1}=-0.0082$ for site 1 and site 2, respectively, which are both close to zero, confirming the validity of the weakly-bound adiabatic approximation, i.e., the decoupling of muon vibrational modes from the vibrational modes of the lattice.
@@
Figure 1: (a) Two crystallographically distinct muon stopping sites in GdRu2Si2, obtained using DFT. (b) Energy barriers between muon sites. Red and blue lines indicate the total (classical plus quantum, zero-point) energy of the muon at sites 1 and 2, respectively. (c) Muon-induced displacements of ions as a function of distance from the lowest-energy muon site.
The energy barriers for moving a muon between sites 1 and 2 were calculated using transition state searches, and confirmed using the nudged elastic band (NEB) method [21], yielding the energy profile in Fig. 1(b). However, on taking into account the quantum zero-point energy, a muon at site 2 (the higher-energy state of the two) has sufficient energy to overcome the energy barrier between its classical location and that of site 1. Site 2 is therefore not stable under quantum fluctuations of the muon position.
While the barrier from site 1 to site 2, and hence the barrier between adjacent instances of site 1 (i.e. site 1 and site $1^{\prime}$), is sufficiently high that we do not expect muons at site 1 to be able to leave this site classically, we also need to consider the possibility of quantum mechanical tunneling. By considering the effective one-dimensional potential for the muon along the minimum-energy path between the classical turning points, i.e., where the black and red lines in Fig. 1(b) intersect, we used the WKB approximation to estimate a tunneling rate $\approx$0.05 Hz for muons to move between site 1 and site $1^{\prime}$ [18], which is significantly longer than the timescale over which measurements are made. We therefore conclude that site 1 is the single crystallographically-distinct stable stopping site in this material and we plot the displacements of the nearby ions in the vicinity of this site in Fig. 1(c). The muon does not introduce any significant perturbations to its local environment, with all displacements $<0.1$, and is therefore expected to be a faithful probe of the magnetism in this system. Note that these calculations do not take into account the possible effect of charge density waves, which have been proposed to result from the coupling between itinerant electrons and localized moments in this system [22], and could result in small perturbations to the muon site.
@@ III Results and Discussion
III.1 Angle-dependent Knight shift measurements
To further examine the muon site and its associated hyperfine coupling, we carried out angle-dependent measurements of the muon Knight shift on a single crystal of GdRu2Si2 with a large [100] face using the GPS spectrometer at the Swiss Muon Source, Paul Scherrer Institute (Switzerland). Measurements were made at 55 K, i.e., within the paramagnetic phase. The sample was enclosed in aluminum foil and attached to the end of a sample stick. These measurements were made in a spin-rotated configuration, with the initial muon-spin polarization making an angle $\approx 45^{\circ}$ to the muon beam axis.
The experimental geometry is shown in Fig 2(a). For the purpose of calculating the magnitude of magnetic field at the muon site, we consider the crystallographic frame ($a$, $b$, $c$) of the sample to be fixed. We then define a second frame, the lab frame, such that the applied magnetic field $\bm{B}_{0}$, with magnitude 0.78 T, initially points along $\bm{\hat{z}}$ and rotates anticlockwise about $\bm{\hat{x}}$, with the angle of rotation denoted $\theta$. The two coordinate systems are related using the Euler angles $\alpha$, $\beta$, and $\gamma$, as shown in Fig 2(a). Assuming perfect alignment of the crystal, i.e., $\alpha=\beta=\gamma=0$, we have $\bm{\hat{z}}=\bm{\hat{a}}$ and $\bm{\hat{x}}=\bm{\hat{b}}$. This corresponds to the scenario in which the applied magnetic field is initially along [100] and the rotation axis of the sample stick is along [010]. The use of Euler angles allows us to capture any sample misalignment in our analysis.
At each rotation angle, we consider the components of the muon polarization measured in the four detectors perpendicular to the applied magnetic field: up, down, left, and right. These are all fit to a function
$A(t)=A_{\mathrm{bg}}\cos(\omega_{0}t+\phi)+\sum_{i=1}^{2}A_{i}e^{-\lambda_{i}t}\cos(\omega_{i}t+\phi),$ (1)
where the first term is due to muons stopping outside the sample and precessing about the applied field, while the other two components correspond to muons stopping in the sample. The phase $\phi$ is allowed to vary for each detector, but is the same for all components in the fitting function, as would be expected for the magnetic field distribution arising from a field-polarized state. Note that it was not possible to separately resolve $\omega_{1}$ and $\omega_{2}$ for $\theta=-90^{\circ}$ and hence these were fixed to be equal in the fitting procedure for this angle (as were $\lambda_{1}$ and $\lambda_{2}$).
For muons at stopping at site 1 we would expect to observe two distinct components with frequencies shifted away from the Larmor frequency $\omega_{0}=\gamma_{\mu}B_{0}$ corresponding to the applied field. For Gd moments pointing in a general direction within the $ac$ plane, instances of site 1 lying on the $a$ axis will experience a different dipolar magnetic field to those lying on the $b$ axis. As each of the magnetically distinct instances of site 1 are expected to be occupied with equal probability, we fix $A_{1}=A_{2}$ in Eq. (1). Furthermore, from a fit to the spectrum measured for $\theta=0$, where the frequency shifts are at their largest, we find $(A_{1}+A_{2})/(A_{1}+A_{2}+A_{\mathrm{bg}})=0.95(2)$. The ratio of the amplitudes is fixed at this value for fits at the other rotation angles. For each component, the precession frequency is related to the magnetic field at the muon site through $\omega_{i}=\gamma_{\mu}B_{i}$. Hence, we can compute the Knight shift at each of the muon sites through
$K_{i}=\frac{\omega_{i}-\omega_{0}}{\omega_{0}}.$ (2)
[…]
Figure 2: (a) Schematic of the experimental geometry, indicating the direction of the applied magnetic field $\bm{B}_{0}$ in the frame ($a$,$b$,$c$) of the sample. (b) Two components of the muon Knight shift as a function of the rotation angle $\theta$.
In this experiment, the contributions to the magnetic field at the muon site are as follows
$\bm{B}(\theta)=\bm{B}_{0}+\bm{B}_{\mathrm{dip}}(\theta)+\bm{B}_{\mathrm{Lor}}+\bm{B}_{\mathrm{cont}}+\bm{B}_{\mathrm{demag}}(\theta),$ (3)
where $\bm{B}_{0}$ is the applied magnetic field, $\bm{B}_{\mathrm{dip}}(\theta)$ is the local dipolar field at the muon site, $\bm{B}_{\mathrm{Lor}}$ is the Lorentz field, $\bm{B}_{\mathrm{cont}}$ is the contact-hyperfine field and $\bm{B}_{\mathrm{demag}}(\theta)$ is the demagnetizing field. The Lorentz term $\bm{B}_{\mathrm{Lor}}$ is proportional to the sample magnetization. We assume that the magnetic susceptibility is independent of field direction within the paramagnetic phase. Hence, we assume Gd3+ moments with a magnitude $\mu_{\mathrm{Gd}}$ in the direction of the applied field, where $\mu_{\mathrm{Gd}}$ is allowed to vary in the fitting, but is assumed to have the same value for all orientations of the applied field. Using this magnetic structure, we can calculate $\bm{B}_{\mathrm{Lor}}$, and, using the dipolar tensor at the muon stopping site, the local dipolar field $\bm{B}_{\mathrm{dip}}$ at each field angle.
The contact term $\bm{B}_{\mathrm{cont}}$ was evaluated using the approach implemented in MuESR [23], whereby the hyperfine field is assumed to be isotropic and results from a scalar coupling to the magnetic moments. Each of the nearest-neighbor moments contributes to the total hyperfine field by an amount inversely proportional to the cube of its distance from the muon. This field is then scaled by the hyperfine coupling constant $A$, leading to
$\bm{B}_{\mathrm{cont}}=A\frac{2\mu_{0}}{3}\sum_{i=1}^{N}\frac{r^{-3}_{i}}{\sum_{j=1}^{N}r^{-3}_{j}}\bm{m}_{i},$ (4)
where $i$ (or $j$) denotes each of the $N$ nearest-neighbor Gd3+ ions at the muon site ($N=2$ in our case). Since all Gd moments are parallel to the applied magnetic field, this term results in a constant shift in both components of the Knight shift proportional to the hyperfine coupling $A$. Furthermore, since the two nearest-neighbor Gd atoms are equidistant from the muon site, they contribute equally to the hyperfine field in Eq. (4) and hence the scaling with distance is not important. The demagnetizing field
$\bm{B}_{\mathrm{demag}}=-\mu_{0}N\bm{M},$ (5)
where the demagnetizing tensor $N$ depends on the sample shape. For the sample used in this experiment $N$ is not isotropic, and therefore the demagnetizing field exhibits an angular dependence.
We evaluated the demagnetizing tensor for our sample using an analytical solution obtained for rectangular prisms [24]. We approximated the sample as a rectangular plate with dimensions 5.7(4)mm by 4.36(2)mm, along $b$ and $c$, respectively, and a thickness of 0.84(2)mm. (The greater uncertainty in the length of the crystal along the $b$ axis reflects the fact that the crystal has two uncut rounded edges). Averaging over the sample volume, this yields a diagonal demagnetizing tensor $N$ with eigenvalues $0.717(2),0.123(5)$ and $0.160(4)$.
The results of our fits to the muon Knight shift at the two muon sites are shown by the red and blue lines in Fig. 2(b) and show good agreement with the data. Best fits are obtained with $\alpha=-64(6)^{\circ}$, $\beta=17(2)^{\circ}$ and $\gamma=68(3)^{\circ}$. This set of Euler angles results in an angle of 16 between the sample rotation axis and [010], while the angle between the initial magnetic field direction and [100] is given by $\beta=17(2)^{\circ}$. Such a degree of misalignment is very plausible, given the experimental setup. The fitting yields a value $A=-0.0094(4)$-3 for the hyperfine coupling constant. The fitted Gd moment is $\mu_{\mathrm{Gd}}=0.46(2)\mu_{\mathrm{B}}$. Magnetometry measurements [18] made on this sample at 0.5 T give a moment of $\mu_{\mathrm{Gd}}=0.437(15)\mu_{\mathrm{B}}$ at 55 K. Assuming that the susceptibility is independent of field at this temperature would give a moment $\mu_{\mathrm{Gd}}=0.68(2)\mu_{\mathrm{B}}$ at 0.78 T, which is slightly larger than the moment suggested by this analysis.
[…]
III.2 Transverse-field $\bm{\mu}^{+}$ measurements
@@
Transverse-field $\mu^{+}$ measurements were made on aligned crystals of GdRu2Si2 using the HAL-9500 spectrometer at the Swiss Muon Source. In these measurements, the muon spin precesses about the sum of the external magnetic field, which is oriented perpendicular to the initial muon-spin direction, and any internal magnetic fields present in the sample. Two sets of measurements were made: one with the crystals oriented such that applied field $\bm{B}_{0}$ pointed along [001] and another set with the crystals oriented such that $\bm{B}_{0}$ pointed along the [100] direction. Spectra were measured after zero-field cooling the sample.
@@
We first consider the spectra measured at $T=40$ K, where we can access the field-polarized state. Fourier transform (FT) $\mu^{+}$ spectra measured at different fields are shown in Fig. 3 for each of the orientations. For $\bm{B}_{0}\parallel[001]$ the spectra comprise a peak close to the applied field and a secondary peak at lower fields which becomes more prominent as the field is increased. (We note that the central peak is split in two for all of the TF spectra measured on this material, which could result from a fraction of the muons stopping in parts of the sample where the internal field almost, but not completely, cancels, such as, for example, domain walls or surfaces.) For $\bm{B}_{0}\parallel$ [100] there is an additional peak at fields larger than $B_{0}$ that is absent for the other orientation. We therefore fit the spectra for fields along [001] and [100] to the sum of two or three Lorentzians, respectively, and plot the behavior of the peak centers in Figs. 3(d) and 3(h). We see that the magnitudes of the field shifts at the two peaks initially increases approximately linearly with the applied field, but then begins to level off, suggesting that the magnetization is starting to saturate. We also note the absence of any satellite peaks for $B_{0}=0.5$ T $\parallel$ [001] in Fig. 3(a), with the magnetic phase diagram for this orientation [shown in Fig. 4(i)] having a phase boundary between 0.5 and 1 T at this temperature. This suggests that the low-field, higher-temperature magnetically ordered state gives rise to a broad distribution of magnetic fields at the muon site, relaxing the contribution of muons stopping in the sample out of the TF spectra, a scenario which is consistent with our ZF spectra, discussed in Sec. III.3, which show a large relaxation rate at this temperature. On the other hand, spectra measured at 0.5 and 1.0 T for $\bm{B}_{0}\parallel[100]$, shown in Figs. 3(e) and 3(f), respectively, are expected to occupy the same magnetic phase [as seen in our ac susceptibility measurements shown in Fig. 3(j)] and both show satellite peaks. This suggests the zero-field ordered state survives up to higher fields applied along [001] than for those along [100].
@@
The number of peaks observed for each of the field orientations reflects the symmetry of the muon stopping site. The local magnetic field at the muon site is the sum of the externally applied field and dipolar, Lorentz, demagnetizing and hyperfine fields. For Gd moments along $c$, muons at site 1 will all experience an internal field of $(-0.063-0.140N+7.77A)\mu_{\mathrm{Gd}}$ T along $c$, where the first term is sum of the dipolar and Lorentz fields, the second term is the demagnetizing field, with the demagnetizing factor $N$ depending on the sample shape, and the final term is the hyperfine field, which is proportional to the hyperfine coupling constant $A$, measured in Å-3. All of these contributions to the internal field are proportional to the Gd magnetic moment $\mu_{\mathrm{Gd}}$, which in this expression is measured in Bohr magnetons. On the other hand, for moments along $a$, instances of site 1 lying on the $b$ axis will experience a field $(0.295-0.140N+7.77A)\mu_{\mathrm{Gd}}$ T along $a$, whereas the crystallographically equivalent site lying on the $a$ axis will instead experience a field $(-0.091-0.140N+7.77A)\mu_{\mathrm{Gd}}$ T. For $\bm{B}_{0}\parallel$ [100] the separation of the satellite peaks, $\Delta B$, is therefore related to $\mu_{\mathrm{Gd}}$ by $\Delta B=0.386\mu_{\mathrm{Gd}}$ T. The development of $\mu_{\mathrm{Gd}}$ as a function of applied magnetic field is shown in Fig. 3(i). We have also measured the Gd moment as a function of magnetic field using magnetometry [18] [Fig. 3(i)]. The values of $\mu_{\mathrm{Gd}}$ obtained using these approaches show reasonable agreement, though we note that those obtained from magnetometry are consistently larger and have a slightly different functional dependence with applied field. Taking the sample geometry to be a flat plate with a perpendicular magnetization, we assume $N\approx 1$. (We note the larger sample area obtained for this mosaic of crystals, compared to a single crystals, means that we expect the out of plane component of the demagnetizing tensor to be closer to 1 than was the case of the sample using in our angle dependent Knight shift measurements). Fitting either the positive or negative field shifts for $\bm{B}_{0}\parallel$ [100] to a straight line in $\mu_{\mathrm{Gd}}(B_{0})$ [blue or red lines in Fig. 3(h), respectively] yields $A=-0.0040(9)~{}\mathrm{\AA}^{-3}$, which is similar to the hyperfine coupling obtained from our analysis of angle-dependent muon Knight shift measurements. This negative hyperfine contribution can be explained by the spin-polarization of conduction electrons through the RKKY interaction, which has been predicted to play an important role in stabilizing the skyrmion lattice in GdRu2Si2 [28]. The multiplicity of the peaks for fields along [100] is consistent with our calculated muon site and with our TF measurements made in the paramagnetic phase at 55 K.